### v2

# Entanglement corner dependence in two-dimensional systems: A tensor network perspective

Abstract
In continuous quantum field theories, the entanglement entropy of a subsystem with sharp corners in its boundary exhibits a universal corner-dependent contribution. We study this contribution via the lens of discretized systems, and demonstrate that this corner dependence emerges naturally from the geometric structure of infinite projected entangled pair states (iPEPS) on discrete lattices. Using a rigorous counting argument, we show that the bond dimension of an iPEPS representation exhibits a corner-dependent term that matches the predicted term in gapped continuous systems.
Crucially, we find that this correspondence only emerges when averaging over all possible lattice orientations, revealing a fundamental requirement for properly discretizing continuous systems. Our results provide a geometric understanding of entanglement corner laws and establish a direct connection between analytical field theory predictions and the structure of tensor network representations. We extend our analysis to gauge-invariant systems, where lattice corners crossed by the bipartition boundary contribute an additional corner-dependent term. These findings offer new insights into the relationship between entanglement in continuous and discrete quantum systems.
preprint: This line only printed with preprint option
I introduction
Tensor networks (TNs), specifically projected entangled pair states (PEPSs) [1], are extremely useful in representing a vast range of physical states on discrete lattices. Their translationally-invariant version, infinite PEPSs (iPEPSs) [2, 3, 1], has proven useful in the study of ground states [4, 5, 6, 7, 8, 9, 10, 11], finite temperature behavior [12, 13, 14, 15, 16, 17, 18, 19], and nonequlibrium [20, 21, 22, 9, 23, 24, 25, 26, 27] of various open problems, and they serve as a state-of-the-art numerical tool for the study of systems that suffer from the sign problem [28].
The explicit relation between the TN representation of a state and its entanglement scaling make TNs a useful analytical tool for the study of entanglement in many-body states (see, e.g., Refs.  [29, 30] on the entanglement of gauge-invariant PEPSs). As explained in detail in Sec. II below, the property referred to as bond dimension of a TN-represented state serves as an upper bound to zeroth Rényi entropy of the state, and in turn, also the von Neumann entropy. Thus, observing the bond dimension behavior allows us to probe the entanglement of TN-representable states. Focusing on 2D-iPEPS-representable states, we study here two-dimensional systems with area-law entanglement (as a leading term in the entanglement scaling).
We summarize our results briefly: We study the entanglement of a subsystem $A$ of a two-dimensional system, with a sharp corner $\theta$ in its boundary, as depicted in Fig. 2a. Refs.  [31, 32, 33, 34, 35, 36, 37] show that in continuous field systems, the entanglement of $A$ and its environment is of the form
$S(A)=b_{\text{vol}}|A|+b_{\text{area}}|\partial A|-b_{A}\left(\theta,|A|,|\partial A|\right),$
where $S(A)$ measures the entanglement of $A$ with its environment (defined in Eq. (2)), $|A|,|\partial A|$ are $A$’s bulk (volume) and boundary (area) sizes, respectively, and $b_{\text{vol}},b_{\text{area}}$ are constant parameters. Note that in critical (gapless) systems, the entanglement scales as an area-law, i.e., $b_{\text{vol}}=0$, and the leading term of the entanglement is proportional to $|\partial A|$ [38, 39]. $b_{A}(\theta)$ is a subleading term representing the dependence of the entanglement on $\theta$, and may generally depend on $|A|,|\partial A|$. Refs. [32, 35] calculate it in several cases: In several gapped condensed matter systems such as fractional quantum Hall (FQH) systems, $b_{A}(\theta,|A|,|\partial A|)=b_{A}(\theta)$. The corner-dependent behavior of continuous systems is further studied in Refs. [40, 34, 41].
In our work, we use the iPEPS structure to show rigorously that in an iPEPS-represented state on a square lattice, a $\theta$-dependent term in the bond dimension term emerges, which does not depend on $|A|,|\partial A|$. We show that the behavior of the emergent term approximates the dependence of $b_{A}(\theta)$ in gapped continuous systems, as obtained in Refs. [32, 35]. This is done by a simple counting argument, similar to the one presented in Ref. [42] for free-fermion systems. By doing this, we connect the analytical field theory arguments with the geometric properties of their discretized versions.
The rest of this paper is organized as follows: In Sec. II we present background on the definition and properties of PEPSs, the entanglement measures used in this work, and the behavior of entanglement in PEPS-represented states. We then present the previously-predicted corner-dependent term in Sec. III, and discuss its emergence on iPEPS-represented states. In Sec. III.1 we show, based on PEPS properties, that the corner-dependent term as predicted for gapped systems emerges in iPEPSs. We demonstrate our argument on general systems, as well as gauge-invariant systems in which the entanglement depends on the lattice corners crossed by the boundary [29]. We conclude in Sec. IV.
II Entanglement in iPEPS and iMPS
Figure 1: (a) A TN is composed of tensors, each corresponding to a lattice site. The leg colored in pink of each tensor is referred to as the physical leg, and corresponds to an index varying over a basis of the Hilbert space of a single lattice site, while legs colored in blue are referred to as the virtual legs, corresponding to summed-over indices corresponding to a virtual Hilbert space with dimension $\chi$, referred to as the bond dimension. A two- (or higher) dimensional TN is referred to as a PEPS, and a one-dimensional TN is referred to as an MPS. (b) The number of nonzero RDM eigenvalues is bounded from above by the total bond dimension across the boundary between the two subsystems, here depicted in yellow.
The PEPS  [1] representation of a two-dimensional square lattice state with $n$ sites and local Hilbert space size $d$ is composed of a two-dimensional array of rank-5 tensors, each tensor corresponding to a lattice site. One index of each tensor, referred to as the physical leg, corresponds to the local Hilbert space of the lattice site represented by the tensor, while the other four indices, referred to as the virtual legs, correspond to a virtual Hilbert space of dimension $\chi$ and represent the correlation between the lattice sites. $\chi$ is referred to as the bond dimension of the tensor. When contracting the virtual legs of neighboring tensors as in Fig. 1a, a $d^{\otimes n}$ tensor is obtained, which can be reshaped into a $d^{n}$ state vector, the standard representation of a quantum state. The one-dimensional analogue of this representation is called an MPS [43], as illustrated in Fig. 1a.
Special cases of PEPSs and MPSs are the iPEPS [2, 3, 1] and iMPS [44]: The state is represented by a single PEPS or MPS tensor, repeated over the lattice and resulting in a translationally invariant state over an infinite system. An important property of iPEPSs and iMPSs used in this work is that the bond dimension $\chi$ is constant across the lattice.
TN representations are inherently related to the entanglement of the state, as we show below. We recap the definition of standard entanglement measures for a system in a pure state |$\Ket{\psi}$. They quantify the entanglement between a subsystem $A$ and its complement $B=\overline{A}$. The reduced density matrix (RDM) of $A$ is defined by
$\rho_{A}=\mathrm{Tr}_{B}\Ket{\psi}\Bra{\psi},$ (1)
where $\mathrm{Tr}_{B}[\cdot]$ stands for a partial trace over the degrees of freedom of $B$. The primary measure of entanglement between $A$ and $B$ is the von Neumann entropy,
$S(A)=-\mathrm{Tr}\left[\rho_{A}\log\rho_{A}\right]=-\sum_{\lambda}\lambda\log\lambda,$ (2)
where $\lambda$ are the RDM eigenvalues. Additional, numerically accessible measures, are the Rényi entropies:
$S^{(n)}(A)=-\log\left(\mathrm{Tr}\left[\rho_{A}^{n}\right]\right)=-\log\left(\sum_{\lambda}\lambda^{n}\right).$ (3)
From Eqs. (2), (3) it is apparent that the entanglement is bounded by the log of the number of nonzero RDM eigenvalues $\lambda$ (its Schmidt rank).
We now bound the entanglement between a subsystem $A$ and its complement in a PEPS-represented state, relying on the Schmidt decomposition [43]: Any state vector of the total system may be written in the form
|$\Ket{\psi}_{AB}=\sum_{i=1}^{N}\psi_{i}\Ket{i}_{A}\Ket{i}_{B},$ (4)
where {|$\{\Ket{i}_{A}\}_{i=1}^{N},\{\Ket{i}_{B}\}_{i=1}^{N},$ are orthonormal states of subsystems $A,B$, respectively, not necessarily spanning fully the Hilbert spaces $\mathcal{H}_{A},\mathcal{H}_{B}$. $N$ is referred to as the Schmidt rank, and $\psi_{i}$ are referred to as the Schmidt values of $\psi$. In the basis above, the RDM of $A$ is obtained by
$\rho_{A}=\sum_{i=1}^{N}|\psi_{i}|^{2}\Ket{i}_{A}\Bra{i}_{A},$ (5)
and it is apparent that the Schmidt values squared are the RDM eigenvalues. Therefore, the number of nonzero Schmidt values $N$ is the number of nonzero RDM eigenvalues, bounding the entanglement.
The formalism above decomposes the state into the local Hilbert spaces of $A$ and $B$, similar to the local decomposition of a PEPS state. Contracting the virtual legs correlating sites inside the same subsystem, as demonstrated in Fig. 1c, represents the state in the decomposed Hilbert space and demonstrates the relation of the virtual bond dimension and the entanglement between $A,B$. The Schmidt rank $N$, and in turn, the entanglement, are bounded by the total bond dimension between the tensors representing $A$ and $B$, as illustrated in Fig. 1b. In iPEPS- and iMPS-represented states this results in an area-law scaling of the entanglement, $N\leq\chi^{|\partial A|}$, which is constant in the one-dimensional case. Note that $N$ is only bounded by the bond dimension $\chi^{|\partial A|}$ and is not equal to it, since the iPEPS representation does not necessarily correspond to the Schmidt form.
III Field theory corner-law entanglement on iPEPS
Figure 2: (a) The studied bipartition between a subsystem $A$ and its complement, where the bipartition boundary contains a sharp corner $\theta$. (b) Following Ref. [35], the results are obtained for a round system of radius $r$ with subsystem $A$ being a circular section with central angle $\theta$. The results are assumed to hold for an infinite system, that is, $r\rightarrow\infty$.
In this section, we study the $\theta$-dependent term in the Schmidt rank of $\rho(A)$, which also relates to the entanglement behavior. We use a counting argument to show that such a term behaves similarly to the behavior of $b(\theta)$ in continuous systems, and approaches the same limits in $\theta\rightarrow 0,\theta\rightarrow\pi$.
We start by describing the expected term in continuous fields: Suppose a two-dimensional system is partitioned into subsystems $A,B$, and that the bipartition’s geometry contains a sharp corner of angle $\theta$, as demonstrated in Fig. 2a. Refs. [32, 35] show that the charge fluctuations between $A,B$ behaves as
$S(A)=b_{\text{vol}}|A|+b_{\text{area}}|\partial A|-b_{A}\left(\theta,|A|,|\partial A|\right),$ (6)
where $|A|,|\partial A|$ correspond to the volume and area of $A$, respectively, and $b_{A}$ depends on $\theta$ and possibly other features of $A$ as well. Ref. [35] finds the universal dependence of $b$ on $\theta$:
$b_{A}(\theta)=-(1+(\pi-\theta)\cot(\theta))\cdot F(A),$ (7)
where $F$ depends on the model properties. This dependence is obtained by studying the entanglement in a pure round system, where the bipartition is composed of two edges emanating the center with an angle $\theta$ between them, as depicted in Fig. 2b. The dependence of $b_{A}$ on $\theta$ is assumed to extend to infinite systems.
$F$ is studied in Ref. [35] for various specific models: In fractional quantum Hall (FQH) and additional gapped states, it is shown that $F$ is independent of $|A|,|\partial A|$:
$b_{A}(\theta)=-(1+(\pi-\theta)\cot(\theta))\cdot\text{const.},$. , (8)
while in critical systems, $F$ is found to depend logarithmically on $|\partial A|$:
$b_{A}(\theta)\propto-(1+(\pi-\theta)\cot(\theta))\cdot\log\left(|\partial A|\right).$ (9)
This behavior is rigorously established in Ref. [35] for charge fluctuations, and is anticipated to generalize to entanglement properties as well, supported by the numerical evidence presented in the same reference. Subsequently, Ref. [45] provides an analytical treatment of the corner-angle dependence for general state properties (referred to as charge cumulants), including entanglement measures. Their analysis reveals that for small opening angles $\theta$, the scaling $b_{A}(\theta)\sim 1/\theta$ emerges, which is consistent with our result in Eq. (7).
Below, we show that the behavior in FQH systems, Eq. (8) may be reproduced in iPEPSs in the continuum limit, at least approximately. Our argument relies on a simple calculation of the bond dimension’s dependence on $\theta$, which in turn, bounds the $\theta$-dependent term in the entanglement, as explained in Sec. II above.
III.1 Emergence of an area-independent corner law
Figure 3: (a) The gray grid indicates the discretized lattice, with the system’s degrees of freedom residing on the lattice sites. The virtual legs crossed by the bipartition are accented in pink, their total number is denoted by $n_{\text{legs}}$. (b) Lattice corners crossed by the bipartition are accented in pink, their total number is denoted by $n_{\text{corners}}$. (c) In order to conserve the rotational invariance of the system, we average $n_{\text{legs}},n_{\text{corners}}$ over different orientations of subsystem $A$. The orientation is denoted throughout the paper by $\varphi$.
We show that a corner-dependent term independent of $|A|,|\partial A|$ is recreated in iPEPSs. We then show that the dependence of this term on $\theta$ equals approximately to the continuous field behavior, Eq. (8). We follow Ref. [35] and study the bipartition in Fig. 2b, where the full system is a round system with radius $r$, and $A$ is a circular section with a central angle $\theta$. One may now bound the entanglement $S(A)$ by bounding the total bond dimension crossed by the bipartition, which, in an iPEPS, equals $\chi^{n_{\text{legs}}}$, where $n_{\text{legs}}$ is the number of virtual legs crossed by the boundary of $A$ (See Fig. 3a). Additionally, we follow Ref. [29], which shows that in gauge-invariant systems, one observes a lattice-corner-law behavior in the entanglement, that is, a contribution to entanglement measures which scales as the number of lattice corners in the system (see Fig. 3b). We therefore also consider separably the number of lattice corners crossed by the bipartition’s boundary, $n_{\text{corners}}$. In order to recover the rotational symmetry of continuous systems, we average over all possible orientations of the subsystem $A$ with respect to the lattice, denoted by $\varphi$, as illustrated in Fig. 3c. The position of the vertex of $\theta$ is averaged over a single unit cell of the lattice. The results are then fitted, based on Eq. (7), to
$f(\theta)=\alpha+\beta(\pi-\theta)\cot\theta.$ (10)
In Fig. 4 we show how the $\theta$ dependence from Eq. (7) is recovered for an iPEPS in the limits $\theta\rightarrow 0$ and $\theta\rightarrow\pi$. The result does not depend on the area $|\partial A|$, i.e., $\beta=\text{const}$. This is true both for the number of virtual legs as well as for the number of lattice corners that reside on the boundary.
Figure 4: (a) The number of virtual legs $n_{\text{legs}}$ crossed by the boundary of $A$ when $\theta\rightarrow 0$ The dependence on $\theta$ emerges in vicinity of its apex: The number of pink virtual legs crossed by the boundary are independent of $\theta$ once averaged over different orientations of $A$. In contrast, the green legs are “skipped” due to $\theta$’s sharpness, and may have been crossed by the boundary if $\theta$ was larger. The number of green legs does not depend on the area $|\partial A|\sim r$. It scales roughly as $a/\sin\theta$, where $a$ is the lattice spacing. In the limit $\theta\rightarrow 0$, this behavior matches the $\theta$-dependent term in $f(\theta)$. While this limit may not be taken fully, since $\tan\theta$ is required to be larger than $a/r$, the $\theta$-dependence of $f$ is already observed as $\theta$ approaches its minimal value. A similar analysis may be applied to $n_{\text{corners}}$. (b) In the limit $\theta\rightarrow\pi$, the virtual legs “skipped” by the boundary of $A$ may only stem from the legs closest to the apex, as may be seen in the illustration, which depicts a single lattice unit cell. The positions of the apex for which the leg will be skipped correspond to the blue area, and it it as also required that the orientation of $A$, denoted by $\varphi$ and demonstrated in the figure, will be $\varphi\in[\theta-\pi/2,\pi/2]$. Performing the integration over the area of the blue rectangle and the orientations fulfilling the requirement, we find that the $\theta$-dependent term of $n_{\text{legs}}$ is proportional to $\frac{a^{2}}{16}(1-(\pi-\theta)\cot\theta)$, exactly as in $f(\theta)$. A similar analysis may be applied to $n_{\text{corners}}$.
In addition to the analysis of $f(\theta)$ for $\theta\rightarrow 0$ and $\theta\rightarrow\pi$, we perform a numerical count of $n_{\text{legs}},n_{\text{corners}}$ as a function of $\theta$ for $\theta\in[0,\pi]$, for varying values of $r$. Note that $r\rightarrow\infty$ may be interpreted as taking the lattice size to 0 and approaching the continuum limit. We use exact counting of virtual legs and lattice corners crossed by the bipartition boundary. The relative variation of $\sim 10^{-2}$ in the extracted coefficients $\beta_{\text{legs}}$ and $\beta_{\text{corners}}$ arises from the least-squares fitting procedure used to extract $\beta$ from the discrete counts. The fits to Eq. (10) were performed over the full range $\theta\in[0,\pi]$, with typical error remaining below $10^{-5}$ large $r$, as shown in Fig. 5c.
In Fig. 5 we show, as an example, the dependence of $n_{\text{legs}}$ on $\theta$ for a particular value of $r$, fitted to Eq. (10). We then fit $n_{\text{legs}},n_{\text{corners}}$ to Eq. (10) for various values of $r$. We plot the dependence, or rather the lack thereof, of the extracted $\beta_{\text{legs}},\beta_{\text{corners}}$ on $r$, demonstrating the independence of $b(\theta)$ on $|\partial A|$. The coefficients $\beta_{\text{legs}},\beta_{\text{corners}}$ are constant as a function of the area, up to a relative variation of $\sim 10^{-2}$, which may account for finite-size effects or numerical errors. In order for the dependence in Eq. (8) to be accounted for, $\alpha$ may and is in fact expected to depend on $|\partial A|$.
We may interpret these results from a different point of view: We may take $|\partial A|=\text{const.}$., and consider increasing $r$ to be the result of decreasing the lattice constant $a$, thus a finer graining of the system, with the appropriate normalization. $r\rightarrow\infty$ thus corresponds to the continuum limit (rather than an infinite system). As may be seen in Fig. 5, the fit to Eq. (10) becomes better as $r$ grows larger, i.e., as the continuum limit is approached. All of the above indicates that $b(\theta)$, the corner-dependent term in the entanglement, is independent of $|\partial A|$, and that Eq. (8) is recovered.
We stress that the bond dimension $\chi$ typically grows with the accuracy as the continuum limit is approached, but as long as it does not depend on the size of subsystem $A$, our argument applies. The corner-dependent term of the full bond dimension (and in turn, the number of nonzero RDM eigenvalues) does not depend on the area, as long as the bond dimension $\chi$ is uniform across the lattice.
Figure 5: (a) Numerical counting of $n_{\text{legs}}$ for various values of $\theta$, fitted to the expected $\theta$-dependence in Eq. (10), for a system with radius $r=4$. (b) The extracted coefficient $\beta$ of Eq. (10) as a function of $r$, is found to remain relatively independent of $r$ (up to a relative error of $\sim 10^{-2}$ ), that is, independent of $|\partial A|$, both for the number of virtual legs and for the number of lattice corners crossed by the bipartition boundary. (c) Fitting quality of $n_{\text{legs}},n_{\text{corners}}$ to the expected $\theta$-dependence as in Eq. (10). In all graphs, the results were obtained by averaging over different orientations of $A$ with resolution $2\pi\cdot 10^{-2}$ and different apex positions with resolution $10^{-1}$ in both dimensions.
III.2 Approaching the continuum limit
The counting argument presented above connects the QFT entanglement corner term with the geometric structure of lattice systems. Although this argument can be extended to other lattice types, it consistently relies on averaging over lattice orientations. This necessity suggests that certain features inherent to continuous systems may be wrongfully represented when a fixed lattice orientation is used. In the following, we further examine aspects of our results in relation to continuum systems and their discretization.
We begin by addressing gauge-invariant systems, where the entanglement has been shown to include a term proportional to $n_{\text{corners}}$, the number of lattice corners crossed by the bipartition boundary [29]. As demonstrated above, both $n_{\text{legs}}$ and $n_{\text{corners}}$ exhibit a similar angular dependence on $\theta$, consistent with the prediction in Eq. (10), but only when averaged over lattice orientations. For a fixed lattice orientation $\varphi$, both $b(\theta)$ and the corner-dependent term become strongly dependent on $\varphi$, and thus the meaning becomes questionable in the continuum limit. Both our result here and the entanglement term found in Ref. [29], indicate that faithfully discretizing a continuous system requires averaging over lattice orientations.
Another important consideration is the case of critical systems, where it is shown in Eq. (9) that $b(\theta)\sim\log|\partial A|$. The $\theta$-dependence of the bond dimension derived above is independent of $|\partial A|$, and determined solely by a fixed region near the corner, with a radius proportional to $\theta$, denoted by $r_{\theta}$. Therefore, our counting argument does not account for the $\theta$-dependence observed in critical systems. A corner term with logarithmic scaling in $|\partial A|$ may arise from the values of the singular spectrum themselves, rather than merely from the count of nonzero singular values, as we do here.
Nonetheless, the similarity in the functional $\theta$-dependence between $b(\theta)$ in critical systems and our count-based result may hint at an underlying connection. It is important to note that our setup involves a finite, round system, where the radius is taken to infinity only after the lattice spacing is taken to zero. In critical systems exhibiting long-range correlations, our argument may need to be adjusted to account for the system being much larger than the correlation length. Developing such a modified counting argument could offer insight into the correct way to discretize critical systems.
IV Conclusions and future outlook
We studied the subleading corner-dependent term in the entanglement of two-dimensional systems, and specifically, its representability by iPEPSs. This corner term is defined for continuous field systems, while iPEPSs represent discretized systems, whether they originally reside on a discretized lattice or serve as an approximation of a continuous system.
We show, based on the bond dimension of the iPEPS and its relation to entanglement, that in the continuum limit a corner-dependent and area-independent term emerges, similar to the one predicted in continuous gapped systems. A similar argument shows that such a behavior may also rise from the lattice-corner-dependence of the entanglement in gauge-invariant systems. Our result ties the entanglement behavior in continuous systems with the geometrical behavior of its discretized version in the general case, as was shown in Ref. [42] for the integrated quantum metric of free fermions.
The technical aspects of our result lead to an important practical insight: the behavior observed in the continuous system could not have emerged from a single fixed lattice orientation $\varphi$. It required averaging over all possible orientations to restore the system’s isotropy. Similarly, this term would not appear if the discretized lattice was not translational invariant, e.g., if it were centered with respect to the system’s center. This emphasizes the necessity of such symmetrization procedures when approximating continuous systems by discretization. It would be interesting to try and obtain a similar result for critical systems using discretization on a fractal, obtaining a similar insight for conformal field theories.
In hindsight, the connection between entanglement in continuous systems and the geometry of their discretized counterparts is perhaps not surprising: the definition of entanglement in such systems requires the introduction of an energy cutoff, that is, a discretization of the system [46]. It would be interesting to explore whether other entanglement properties, across different spatial dimensions, can similarly be linked to geometric features. Extending the argument to critical systems with long correlation lengths also presents a compelling direction for future investigation.
Appendix A Bond dimension corner dependence when $\theta\rightarrow\pi$
In this section, we obtain explicitly the dependence of $n_{\text{legs}},n_{\text{corners}}$ on $\theta$ for $\theta\rightarrow\pi$, and show that the dependence matches Eq. (7) as in QFT.
Recall that we average $n_{\text{legs}},n_{\text{corners}}$ over the orientation of subsystem $A$ denoted by $\varphi$, and over the vertex position in the unit cell. We start with the dependence of $n_{\text{legs}}$ on $\theta$. When $\theta\rightarrow\pi$, its effect on $n_{\text{legs}}$ stems from lattice sites closest to the vertex, which may be skipped or picked up by the boundary, depending on $\theta$, see Fig. 6a. For a given corner $\theta$ and orientation $\varphi$ and varying vertex position, the probability of skipping the nearest lattice sites (see the geometric depiction in Fig. 6b):
$p_{\text{skip}}(\theta,\varphi)=\frac{1}{2(\tan(\theta-\varphi)+\tan\varphi)}.$ (11)
We now integrate over the orientation $\varphi$. Note that the expression above is positive for $\varphi-m\pi/2\in(\theta-\pi/2,\pi/2)$ where $m$ is an integer. For orientations outside of this range, $p_{\text{skip}}$ may be geometrically shown to be 0, as demonstrated in Fig. 6c. The final expression is obtained to be
[…]
with a corner dependence identical to the continuous system behavior in Eq. (7). The above analysis applies for all $\theta$s large enough such that they will surely pick up the bonds in the neighboring plaquette, which in a square lattice requires $\theta<2\tan^{-1}(1/2)\approx 0.3\pi$. Therefore, Eq. (12) applies to the range $0.3\pi<\theta<\pi$, and not only $\theta\rightarrow\pi$.
We now focus on $n_{\text{corners}}$. In Fig. 6d, we depict the vertex positions in which a boundary corner $\theta$ of orientation $\varphi$ picks up a corner of the lattice. Here we consider different ranges of orientations $\varphi$. The probability to pick up a corner when $\varphi\in(\theta-\pi/2,\pi/2)$ is again obtained geometrically:
$p_{\text{corner}}(\theta,\varphi)=\frac{1}{2}\cot\varphi+\frac{1}{2}\cot(\theta-\varphi)-2p_{\text{skip}}(\theta,\varphi).$ (13)
Integrating over the orientations, one obtains
$p_{\text{corner}}(\theta)=\int_{\theta-\pi/2}^{\pi/2}p_{\text{corner}}=(\theta,\varphi)d\varphi=\log\cos(\pi-\theta)-2p_{\text{skip}}(\theta).$ (14)
Since $\theta\rightarrow\pi$, we obtain
$p_{\text{corner}}(\theta)=-2p_{\text{skip}}(\theta)+O\left((\pi-\theta)^{2}\right),$ (15)
obtaining the behavior of Eq. (7), similarly to $n_{\text{legs}}$. Equivalent calculations may be done for the remaining values of $\varphi$ to obtain a similar dependence on $\pi-\theta$.
Figure 6: (a) When $\theta\rightarrow\pi$, its effect on $n_{\text{legs}}$ stems from the bond legs closest to the vertex. We observe a corner $\theta$ with orientation $\varphi$. Different positions in the vertex result in different contributions to $n_{\text{legs}}$. The red edges crosses the top and bottom bonds, while the blue edges do not pick up any lattice site in the unit cell, i.e., no lattice site is included in $A$. Thus, the edges do not cross any bond. (b) For given $\theta,\varphi$, the blue area denotes the vertex positions which skip a bond. It may be geometrically shown to be $a^{2}/2(\tan\left(\theta-\varphi)+\tan\varphi\right)$. (c) When $\varphi<\theta-\pi/2$, it is geometrically apparent that at least on of the lattice sites is included in $A$ (see the site colored in red). Thus, $p_{\text{skip}}(\theta,\varphi)=0$ in these orientations. (d) For given $\theta,\varphi$, the unit cell subsets which contribute to $n_{\text{corners}}$ are colored in light blue. Example edges are added in dark blue for each of the cases.